OpenStudy (anonymous):

y=10 is how much you get a week for your allowance and x is how much you get payed a week her is my table x(week) y(money)( 0,0)( 1 ,10)( 2 ,20) (3, 30) what i need help on is this Write the function rule based on your table of values. Graph each of your functions on the coordinate plane. can u guys show me how on this first one so i could do the other one on my own thanks so much

jimthompson5910 (jim_thompson5910):

When x = 0, y is y = 0 When x = 1, y is y = 10 So the slope is (change in y)/(change in x) = (10-0)/(1-0) = 10 So y = mx+b becomes y = 10x+b We know that when x = 0, y = 0, so plug these values in to get y =10x+b 0 = 10(0) + b 0 = 0 + b 0 = b b = 0 So the function is y = 10x which you can write as f(x) = 10x

OpenStudy (anonymous):

so what does mx+b

OpenStudy (anonymous):

so from there i dont get it

jimthompson5910 (jim_thompson5910):

y = 10x is in the form y = mx+b where m = 10 and b = 0

jimthompson5910 (jim_thompson5910):

Does that make sense?

OpenStudy (anonymous):

what is the b

jimthompson5910 (jim_thompson5910):

b = 0, which is the y-intercept

jimthompson5910 (jim_thompson5910):

So this means that the y-intercept is (0,0)

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

so f(x) = 10x is that the answer to this question

jimthompson5910 (jim_thompson5910):

yes, that's the function that represents this idea

OpenStudy (anonymous):

lol thanks

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

also do u know how to graph it could u show me how to do it

jimthompson5910 (jim_thompson5910):

do you know how to plot points in a coordinate system?

OpenStudy (anonymous):

ya a little

jimthompson5910 (jim_thompson5910):

To graph, you just plot two points, say (0,0) and (1,10). Then you draw a straight line through them to graph the line.

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):